OpenStudy (anonymous):

3z-30=5x -7x-x help thank you ^^

OpenStudy (anonymous):

do you mean 3x? or is it 3z

OpenStudy (anonymous):

aww yeah 3x sorry

OpenStudy (anonymous):

thats cool just checkin' 3x - 30 = 5x - 7x -x when you have something like this, with one variable (x) and some number (in this case 30) you want to get the variable (x) to one side, and the other numbers on the other side for instance: 3x - 30 = 5x - 7x -x 3x - 5x + 7x + x = 30 you can simplify further by adding/subtracting the x's 8x + 3x - 5x = 30 or 6x = 30 and then find x

OpenStudy (anonymous):

oh thank you, i'll work through it see what happens.